OpenStudy (anonymous):

Check please Identify the vertical asymptote(s) of the function f(x) =x^2-1/2x^2-8

OpenStudy (anonymous):

would it be x=0?

OpenStudy (anonymous):

@tcarroll010 can you check my answer

OpenStudy (anonymous):

The vertical asymptote is where you have unallowable values in the denominator. So, if your full denominator is (x^2 - 8). Is that your denominator? If it is, then see where that expression equals "0".

OpenStudy (anonymous):

it would be 2

OpenStudy (anonymous):

Otherwise, if the expression is: x^2 - 1/(2x^2) - 8 then you are correct.

OpenStudy (anonymous):

The reason I am presenting you with 2 choices is that I can't tell if that expression is really supposed to be the one in my last post or if you mean: (x^2 - 1) / (2x^2 - 8)

OpenStudy (anonymous):

without parenthesis ill put a pic so i can know where to put the parenthesis next time

OpenStudy (anonymous):

Okay, so your denominator is : (2x^2 - 8) and that will be 0 at 2 different values of x. Just solve: 2x^2 - 8 = 0

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

That's one of the values. Don't forget that you are dealing with a square root. Do you remember something about positive and negative roots?

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

Hint: sqrt(x^2) = +-x

OpenStudy (anonymous):

x=2 x=-2

OpenStudy (anonymous):

horizontal is y=1/2

OpenStudy (anonymous):

Yes! You got It!